OpenStudy (anonymous):

Solve the exponential equation 3^x^2+5x=1/81

OpenStudy (anonymous):

\[\large 3^{x^2+5x}=\frac{1}{81}\] right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

since \(3^{-4}=\frac{1}{3^3}=\frac{1}{81}\) you know that \[x^2+5x=-4\] solve that quadratic

OpenStudy (anonymous):

more explain plz

OpenStudy (anonymous):

(x-5)(x+1) x=5 x=-1